How many moles of hydrogen chloride are contained in a 52.6-g sample of this gas?
1
Expert's answer
2020-04-08T13:27:55-0400

The number of moles can be determined if we divide the mass of the sample by its molar mass:


"n=\frac{m}{\Mu}=\frac{52.6}{1+35.5}=1.44\text{ mol}."